OpenStudy (anonymous):

lim x→13 (x^2−19x+78)/(x−13.)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 13} \frac{ x²-19x+78 }{ x-13 }\]

OpenStudy (anonymous):

Ok I guess I'm supposed to factorise the numerator?

OpenStudy (harsimran_hs4):

yes you are on the right track ....so can you factorize it?

OpenStudy (anonymous):

uh.. 78/13=6 but (x-13)(x+6)=x²-12x+78

OpenStudy (anonymous):

so it could be (x-13)(x+6)+7x?

OpenStudy (harsimran_hs4):

(x-13) (x-6) is the required factorization

OpenStudy (anonymous):

ok so if it would be (x-13)(x-6)+7x, then (x-13) cancels out and we're left with x-6+7x=6+6x, when we plug in x=13, we get 6+78=84 ?

OpenStudy (harsimran_hs4):

what is 7x doing there?

OpenStudy (harsimran_hs4):

(x-13) (x-6) /(x-13)= (x-6) 13-6 = 7 so 7 is the answer

OpenStudy (anonymous):

all righty now I got it... the seven was a huge mistake that's what it was

OpenStudy (harsimran_hs4):

cool :)